OpenStudy (anonymous):

Prove: tanx+secx=cosx/1-sinx

OpenStudy (anonymous):

tanx = (sinx)/(cosx) secx = 1/(cosx) ⇒ tanx +secx = (sinx)/(cosx) + 1/(cosx) ⇒ tanx + secx = (sinx + 1)/(cosx) ⇒(sinx + 1)/(cosx) = (cosx)/(1 - sinx) is equivalent to: (sinx + 1).(1 - sinx) = (cosx).(cosx) = = (1 + sinx).(1 - sinx) = (cosx)² ⇒ (1 + sinx).(1 – sinx) = 1 – sin² x = cos² x

OpenStudy (anonymous):

that doesnt seem right. I dont know if you understood the problem right

OpenStudy (anonymous):

@andrea408

OpenStudy (anonymous):

cosx/(1-sinx) X (1+sinx)/(1+sinx) cosx(1+sinx)/(1-sin^2x) cosx(1+sinx)/(cos^2x) (1+sinx)/(cosx) 1/cosx + sinx/cosx secx + tanx

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@andrea408 what side are you suppose to start from?

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/1991/prove?show=1992#a1992 ask your trigonometry homework questions at http://www.mathskey.com/question2answer/ … and get free math help. all the best